OpenStudy (anonymous):

Number of Days: Total Cost of Rent: 1 $54 3 $112 5 $170 7 $228 Is this a linear relation?

OpenStudy (anonymous):

(1,54) and (3,112) (112-54)/(3-1)=58/2=29 (5,170) and (7,228) (228-170)/(7-5)=58/2=29 so yes they are linear, hope this helped